Please help me solve:
Solve the equation. Check the solution.
-2 over x+4= 4 over x+3

Answer by jorel1380(2518)

-2/x+4=4/x+3
-2(x+3)=4(x+4)
-2x-6=4x+16
6x=-22
x=-22/6..